# Gravitational waves from a dilaton-induced, first-order QCD phase transition

### v2

@@ Abstract
In this paper, we show that the dynamics of $\phi$ can effectively turn the QCD phase transition first-order, leading to detonating bubbles of the confined phase, separated from the quark-gluon plasma by a dilaton domain wall. The expanding bubbles quickly reach relativistic speeds and produce a stochastic gravitational wave background with an amplitude and spectrum that agree well with recent observations from Pulsar Timing Arrays [4, 5, 6, 7], thus begging the question: Have we already observed the first signals from a dilaton-induced first-order QCD phase transition?
@@
We assume that the zero-temperature dilaton potential, $V_{0}(\phi)$, has two minima: the true vacuum at $\phi_{\rm TV}$ and the false vacuum at $\phi_{\rm FV}$, with $\Delta V=V_{0}(\phi_{\rm FV})-V_{0}(\phi_{\rm TV})>0$, and, importantly, $\Delta\mathfrak{f}=\mathfrak{f}(\phi_{\rm FV})-\mathfrak{f}(\phi_{\rm TV})>0$. This assumption means that QCD is more weakly coupled in the false vacuum than in the true vacuum. We do not model the microphysical origins of $V_{0}(\phi)$ and $\mathfrak{f}(\phi)$. In potential string theory realisations of our scenario, $\phi$ would be a modulus field whose potential can be quite complex, receiving both perturbative and non-perturbative contributions [8, 9, 10]. Similarly, the gauge kinetic function may descend from the holomorphic gauge kinetic function of supergravity, which receives both one-loop and non-perturbative corrections [11, 12]. We provide a concrete toy model realising our scenario below, however, many generalisations are possible. Recent work on related theories include [13, 14, 15, 16, 17, 18, 19, 20].
In the present universe, $\phi=\phi_{\rm TV}$, and the properties of the dilaton are experimentally constrained. The dilaton mass is required to be sufficiently large to avoid collider constraints [21]: $m_{\phi\,{\rm TV}}\gtrsim 5\,{\rm TeV}$, under some assumptions on the form of $\mathfrak{f}$. In this work, we assume $m_{\phi\,{\rm TV}}\sim m_{\phi\,{\rm FV}}\sim 10\,{\rm TeV}$, for simplicity. Moreover, the QCD coupling constant is set by experiments to $\alpha_{s}(m^{2}_{Z})=0.1180\pm 0.0009$ (cf. [22] and references therein), which implies
@@
denotes the one-loop thermal effective potential due to dilaton self-interactions [27]. Here $J_{B}$ is the thermal bosonic function and $m_{\phi}(\phi)$ is the field-dependent mass. The term $V_{\cal P}(\phi)$ represents the thermal QCD pressure-corrections ($\mathcal{P}_{\rm QCD}=-V_{\cal P}$) in a fixed dilaton background [28]:
$V_{\cal P}(\phi)=-\frac{8\pi^{2}T^{4}}{45}\left(\frac{17}{3}-\frac{235}{16}\alpha_{s}(\phi)+\mathcal{O}(\alpha_{s}^{3/2})\right)\;,$ (7)
where $\alpha_{s}(\phi)=1/(4\pi\mathfrak{f}(\phi))$ 222 This correction is valid for the $\phi$ values we are interested in at temperatures higher than the confinement scale, since the RG-evolved coupling satisfies $\alpha_{s}(\phi_{\rm TV},\mu=T)<1$, so that QCD is still perturbative.,  [^3] We note that the interactions of dilaton fluctuations with the gluons with vertex structure $\sim(p_{1}^{\mu}p_{2}^{\nu}-g^{\mu\nu})$ renormalise the couplings but do not have finite contributions at $T\neq 0$ that could change the shape of the 1-loop thermal potential, and hence we neglect these contributions..
@@
At intermediate temperatures, the thermal contributions from self-interactions become Boltzmann-suppressed and can be ignored, and $V_{\rm eff}$ develops a second minimum at smaller field values. Notably, due to $V_{\cal P}$, the large-field minimum approaching $\phi_{\rm FV}$ is the global minimum of $V_{\rm eff}$ at these temperatures, and the dilaton is attracted to the false vacuum.
@@
Figure 1: Left: The dilaton potential, $V_{\rm eff}(\phi)$, at different temperatures. The values of the potential are rescaled for visual clarity. For $T\sim 100~{}\mathrm{TeV}\gg m_{\phi}$ the minimum is located in between the tree-level vacua. As temperature drops ($T\sim\mathcal{O}(10)~{}\mathrm{GeV}\ll m_{\phi}$), the minimum moves towards the $\phi_{\rm FV}$ while a local minimum emerges near $\phi_{\rm TV}$. When temperature corrections become unimportant, $\phi_{\rm TV}$ becomes the global minimum. Right: Schematic contour plot for the two-field potential, $V_{\rm tot}(\chi,\phi)$ for $T_{\rm QCD}^{\rm FV}<T<T_{\rm QCD}^{\rm TV}$.
@@
With $\phi$ in the false vacuum, the quark-gluon plasma may cool below $T_{\rm QCD}^{\rm TV}$ without triggering the cross-over transition to the hadron gas (which for $\phi=\phi_{\rm FV}$ would happen around $T_{\rm QCD}^{\rm FV}$). However, the dilaton vacuum is non-perturbatively unstable and eventually decays to $\phi_{\rm TV}$ through quantum tunnelling, proceeding through the nucleation of true-vacuum bubbles. The nucleation rate per unit volume is given by $\Gamma/\mathcal{V}=Ae^{-S_{4}}$ [31, 32], with $S_{4}$ the time independent $O(4)$ symmetric bounce action, and $A$ a prefactor. We assume that the nucleation temperature, $T_{n}$, is in the range
$T_{\rm QCD}^{\rm FV}<T_{n}<T_{\rm QCD}^{\rm TV}\,,$ (8)
and that $T_{n}\gtrsim T_{\rm BBN}\approx 6\;{\rm MeV}$ [33]. Fixing the nucleation temperature corresponds to fixing the decay rate through the approximate equation $\Gamma/{\cal V}\simeq H^{4}(T_{n})$, expressing that, at $T_{n}$, one bubble is nucleated per Hubble volume and Hubble time.
@@
Outside the bubbles, the plasma remains in the pre-confined quark-gluon phase (as $T_{n}>T_{\rm QCD}^{\rm FV}$), but inside the bubbles, the temperature is lower than the local QCD confinement scale (i.e. $T_{\rm QCD}^{\rm TV}$), leading to prompt hadronisation and chiral symmetry breaking. In the two quark linear sigma model, the approximate timescale of confinement can be found from $t_{\rm roll}\sim{\cal O}(1/m_{\chi})\ll H^{-1}$, with $m_{\chi}=\sqrt{2\lambda f_{\pi}}$, so chiral-symmetry breaking inside the true-dilaton-vacuum bubbles is approximately instantaneous. Thus, in this scenario, the QCD PT is no longer a smooth cross-over, but characterised by expanding bubbles of the hadronic phase inside the quark-gluon plasma; this scenario effectively turns the QCD phase transition first order.
@@
We require that the transition percolates (at plasma temperature $T_{\star}$) which, for a PT with constant $\Gamma$, requires $\alpha\lesssim 20$ [34], where $\alpha=\Delta V/\rho_{\rm rad}(T_{\rm n})$. For such transitions, the inverse duration of the PT, $\beta$, is set by $\alpha$ and satisfies $38\gtrsim\beta/H_{*}>$ 3 [34]. As the energy density $\Delta V$ is channelled into the Standard Model, the plasma is reheated to the temperature $T_{\rm rh}=(1+\alpha)^{1/4}T_{n}$, here assuming $T_{\star}\approx T_{n}$. We restrict our discussion to $T_{\rm rh}\lesssim T_{\rm QCD}^{\rm TV}$, so that the hadrons do not deconfine (and then undergo a ‘second’ QCD transition, this time a cross-over).
@@
The subsequent cosmology and observational predictions of this scenario depend on the speed at which the bubbles expand. As we show in the Supplementary Material [35], the bubble walls reach ultrarelativistic terminal velocities before bubbles collide.
@@
Gravitational wave signal — As the relativistic dilaton bubbles grow large, collide, and percolate, they produce a stochastic gravitational wave (GW) background, potentially observable today [36, 37]. This is particularly interesting given the recent observation of the Hellings-Down correlation in Pulsar Timing Array (PTA) datasets [4, 5, 6, 7], providing evidence for a stochastic GW background in the nanohertz range. Such a background could potentially be explained astrophysically through supermassive black hole binary mergers [38], but this explanation has difficulties with accommodating the observed amplitude and spectrum [39, 40]. By contrast, several studies have found that GWs generated by a first-order PT provides a better fit to the data [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (see [56] for an alternative gravitational wave explanation generated by turbulent sources regardless of the order of the QCD PT).
Figure 2: Gravitational wave spectra from a dilaton-induced QCD PT overlaid with NANOGrav 15-year free-spectrum constraints (gray violins). Left: Spectra for transitions at fixed temperature $100\,\rm MeV$ with varying strength, $\alpha=0.7$, $3$, $6.8$. Right: Spectra for fixed strength $\alpha=6.8$ and varying temperature, $T_{n}=50$, $75$, $100~{}{\rm MeV}$.
@@
Since the bubbles are expected to reach a terminal velocity before they collide, the main contribution to the GW signal in this scenario is from the dynamics of sound waves generated in the plasma [57, 58, 59, 60, 61] (the contribution to the signal coming from bubble wall collisions, which is relevant for runaway PTs, is discussed in the Supplementary Material [35]).
@@
[…]
@@
×$\times\left(\frac{61.75}{g_{*}(T_{\star})}\right)^{1/3}v_{w}\left(\frac{f}{f_{\rm sw}}\right)^{3}\left(\frac{7}{4+3\left(f/f{\rm sw}\right)^{2}}\right)^{7/2}\;,$ (10)
@@
In Fig. 2 we show the GW energy spectrum generated by the dilaton-induced phase transition for several benchmark values of the parameters. In the right panel we vary the transition strength ($\alpha=0.7$, $3$ and $6$) and set $T_{n}=100~{}{\rm MeV}$, while in the left panel the temperature is varied ($T_{n}=50~{}{\rm MeV}$, $75~{}{\rm MeV}$ and $100~{}{\rm MeV}$) and $\alpha=6.8$. The spectra are overlaid with the posterior distributions from the NANOGrav 15-year free-spectrum analysis [4], shown as gray violins. Remarkably, the predictions of our model are broadly compatible with the observed spectrum.
@@
To realise the hierarchy of Eq. (8) with a nucleation temperature of $T_{n}=100\,{\rm MeV}$ one may, for example, choose $m\sim 10\,{\rm TeV}$, $\Delta\phi\simeq 2\times 10^{-3}~{}{\rm GeV}$, $\Delta\phi/\Lambda_{1}\sim\mathcal{O}(1)$ and $\Delta V\simeq 0.7\rho_{\rm rad}(T=100\text{MeV})$. In this case the barrier height in $\phi$ direction is large, and the thin-wall approximation can be used to estimate $S_{4}$ and the decay rate. Thus, our scenario can be realised with a rather minimal set of ingredients.
Discussion — We have shown that quantum tunnelling of the QCD dilaton can turn the QCD phase transition first order and generate a gravitational wave signal intriguingly similar to that observed by Pulsar Timing Arrays.
@@
Finally, while the explicit model presented in this work illustrates how it may be realised with rather simple ingredients, it would be interesting to embed this toy model in a more complete theoretical framework.
Acknowledgments—We thank Mark Hindmarsh and Filippo Sala for stimulating discussions. This work was supported by the Swedish Research Council (VR) under grants 2018-03641, 2019-02337 and 2024-04289. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@
In this section, we calculate the pressure exerted on the wall by the quark gluon plasma and determine the terminal speed of the bubble walls. Recall that we consider percolating phase transitions with time-independent $\Gamma$. The $\beta$-parameter is given by $\beta=\dot{I}(t_{\star})$, with $I(t)$ the average number of bubbles in the past comoving light cone [34], i.e.
[…]
where $a(t)$ denotes the scale factor, $r_{\rm com}(t_{\star},t)$ the comoving bubble radius and $t_{\star}$ the percolation time, defined by $I(t_{\star})=1$. As mentioned in the main text, for transitions that percolate, $0<\alpha\lesssim 20$, which corresponds to $38\gtrsim\beta/H_{*}>$ 3
@@
where $g_{q}^{*},\,g^{*}_{g}$ denote, respectively, the effective degrees of freedom for quarks and gluons in the plasma. The inequality assumes $\alpha\gtrsim 3\cdot 10^{-3}\,(\Lambda_{\rm QCD}/T_{n})^{2}$, which holds across the most relevant parameter space. Thus, one-to-one processes alone do not suffice to stop the runaway behaviour [89] 444 A more accurate description of friction in a confinement phase transition is given in [90], and results in the scaling $\mathcal{P}_{\rm friction}=\mathcal{P}_{\rm BS}\sim\sum_{i\in BS}\frac{m_{i}^{2}T_{n}^{2}}{24}$, which coincides with that of [89], used here. .
@@
with $\phi_{w}$ a value of the dilaton on the wall and $N_{g}=8$ is the number of gluons. Here, we have assumed that QCD remains perturbative at the wall, i.e. for an intermediate value of $\phi$. This is typically justified since $\Lambda_{\rm QCD}(\phi_{w})\ll\Lambda^{\rm TV}_{\rm QCD}$. The pressure contribution $\mathcal{P}_{1\rightarrow 2}$ grows with $\gamma$ and eventually catches up with $\Delta V$, leading to an ultra-relativistic terminal velocity with gamma factor  [^5] Another potential friction source is bound state formation from string fragmentation [77]. However, due to insignificant supercooling, we expect confinement to happen among quark pairs rather than quarks forming flux-tubes attached to the bubble wall, and string fragmentation to be unimportant. Even in this scenario though, the leading and next-to-leading order friction from gluons has similar dependence on $\gamma$
@@
Figure S1: Gravitational wave spectra from a dilaton-induced QCD PT overlaid with NANOGrav 15-year free-spectrum constraints (gray violins), including the contribution from bubble wall collisions. Solid lines show the spectra in the regime where the energy is carried by bubble walls, while dashed lines correspond to the regime where the energy is converted into sound waves, for comparison. Left: Spectra for transitions at fixed temperature $100\,\rm MeV$ with varying strength, $\alpha=0.7$, $3$, $6.8$. Right: Spectra for fixed strength $\alpha=6.8$ and varying temperature, $T_{n}=50$, $75$, $100~{}{\rm MeV}$.

### v3

@@ Abstract
In this paper, we show that the dynamics of $\phi$ can effectively turn the QCD phase transition first-order, leading to detonating bubbles of the confined phase, separated from the quark-gluon plasma by a dilaton domain wall. The expanding bubbles quickly reach relativistic speeds and produce a stochastic gravitational wave background with an amplitude and spectrum that agree well with recent observations from Pulsar Timing Arrays (PTAs) [4, 5, 6, 7], thus begging the question: Have we already observed the first signals from a dilaton-induced first-order QCD phase transition?
The mechanism introduced in this Letter is based on multiple-field dynamics and quantum vacuum transitions, and should be contrasted to the much-studied scenario of thermal PTs. It is well established that physics Beyond the Standard Model (BSM) can turn the QCD cross-over into a thermal first-order PT. If the resulting PT is sufficiently strong, such models can produce gravitational waves in the frequency band probed by PTAs [8, 9, 10]. However, recent work has shown that the relevant parameter space is hard to realise in explicit models, either due to weak transition strengths [11, 12, 13], non-relativistic wall velocities [14], or failure to percolate [15]. Conversely, we will now show that the dilaton-induced QCD PT naturally leads to relativistic bubble walls and offers a compelling, and rather minimal, explanation of the PTA observations.
@@
We assume that the zero-temperature dilaton potential, $V_{0}(\phi)$, has two minima: the true vacuum at $\phi_{\rm TV}$ and the false vacuum at $\phi_{\rm FV}$, with $\Delta V=V_{0}(\phi_{\rm FV})-V_{0}(\phi_{\rm TV})>0$, and, importantly, $\Delta\mathfrak{f}=\mathfrak{f}(\phi_{\rm FV})-\mathfrak{f}(\phi_{\rm TV})>0$. This assumption means that QCD is more weakly coupled in the false vacuum than in the true vacuum. We do not model the microphysical origins of $V_{0}(\phi)$ and $\mathfrak{f}(\phi)$, but note that the requisite structure can be readily realised with simple ingredients. Scalar potentials with localised bumps (or dips) can, e.g., be generated through the Coleman-Weinberg potential by integrating out bosons (or fermions) whose masses are minimised at finite values of $\phi$. In potential string theory realisations of our scenario, $\phi$ would be a modulus field whose potential can be quite complex, receiving both perturbative and non-perturbative contributions [16, 17, 18]. Stabilised string compactifications with positive vacuum energy automatically involve multiple local minima (some asymptotic in field space) [19, 20], and long-standing general arguments suggest that our universe is located in regions of field space where the scalar potential is rather complex, with multiple important contributions [21, 22], generating a rich landscape of local minima. Similarly, the gauge kinetic function may descend from the holomorphic gauge kinetic function of supergravity, which receives both one-loop and non-perturbative corrections [23, 24]. Below, we provide a concrete toy model that realises our scenario and draws on ingredients commonly found in string compactifications, however, many generalisations are possible. Recent work on related theories include [25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
In the present universe, $\phi=\phi_{\rm TV}$, and the properties of the dilaton are experimentally constrained. Collider searches at the Large Hadron Collider (LHC) have been used to constrain dilaton theories with linear gauge kinetic functions, $f\sim\phi/\Lambda$, leading to limits on $\Lambda$ in the TeV-range for dilaton masses $2\,{\rm TeV}\leq m_{\phi}\leq 5\,{\rm TeV}$ [35]. In this work, we assume $m_{\phi\,{\rm TV}}\sim m_{\phi\,{\rm FV}}\sim 10\,{\rm TeV}$, for simplicity. Moreover, the QCD coupling constant is set by experiments to $\alpha_{s}(m^{2}_{Z})=0.1180\pm 0.0009$ (cf. [36] and references therein), which implies
@@
denotes the one-loop thermal effective potential due to dilaton self-interactions [41]. Here, $J_{B}$ is the thermal bosonic function and $m_{\phi}(\phi)$ is the field-dependent mass. The term $V_{\cal P}(\phi)$ represents the thermal QCD pressure-corrections ($\mathcal{P}_{\rm QCD}=-V_{\cal P}$) in a fixed dilaton background for $N_{c}=3$ and $N_{f}$ flavours [42]:
$V_{\cal P}(\phi)=$ $-\frac{8\pi^{2}}{45}T^{4}\Bigg(\left(1+\frac{21N_{f}}{32}\right)$
$-\frac{15\alpha_{s}(\phi,T)}{4\pi}\left(1+\frac{5N_{f}}{12}\right)+\mathcal{O}(\alpha_{s}^{3/2})\Bigg)\;\;,$ (7)
where $\alpha_{s}(\phi)=1/(4\pi\mathfrak{f}(\phi))$. The contribution of the pressure term due to the coupling of the dilaton to the Standard Model in the context of PTs was discussed recently in [43] and in the context of dark matter in [44]. [^2] This correction is valid for the $\phi$ values we are interested in at temperatures higher than the confinement scale, since the RG-evolved coupling satisfies $\alpha_{s}(\phi_{\rm TV},\mu=T)<1$, so that QCD is still perturbative. For temperatures where QCD is non-perturbative near $\phi_{\rm TV}$, we use the fit to lattice calculations from [50] (c.f. the Supplemental Material [52] for a more detailed discussion). [^3] We note that the interactions of dilaton fluctuations with the gluons with vertex structure $\sim(p_{1}^{\mu}p_{2}^{\nu}-g^{\mu\nu})$ renormalise the couplings but do not have finite contributions at $T\neq 0$ that could change the shape of the 1-loop thermal potential, and hence we neglect these contributions..
@@
At intermediate temperatures, the thermal contributions from self-interactions become Boltzmann-suppressed and can be ignored, and $V_{\rm eff}$ develops a second minimum at smaller field values. Notably, due to $V_{\cal P}$, the large-field minimum approaching $\phi_{\rm FV}$ is the global minimum of $V_{\rm eff}$ at these temperatures, and the dilaton is attracted to the false vacuum. This is a key property of our scenario and a general feature of theories with field-dependent couplings: the theory is pushed to weaker couplings at high temperatures. This effect traps the dilaton at $\phi_{\rm FV}$, providing natural initial conditions for the weak-to-strong PT.
@@
Figure 1: Left: The dilaton potential, $V_{\rm eff}(\phi)$, at different temperatures. The values of the potential are rescaled for visual clarity. For $T\sim 100~\mathrm{TeV}\gg m_{\phi}$ the minimum is located in between the tree-level vacua. As temperature drops ($T\sim\mathcal{O}({\color[rgb]{0,0,1}1})~\mathrm{GeV}\ll m_{\phi}$), the minimum moves towards the $\phi_{\rm FV}$ while a local minimum emerges near $\phi_{\rm TV}$. When temperature corrections become unimportant, $\phi_{\rm TV}$ becomes the global minimum. Right: Schematic contour plot for the two-field potential, $V_{\rm tot}(\chi,\phi)$ for $T_{\rm QCD}^{\rm FV}<T<T_{\rm QCD}^{\rm TV}$.
@@
With $\phi$ in the false vacuum, the quark-gluon plasma may cool below $T_{\rm QCD}^{\rm TV}$ without triggering the cross-over transition to the hadron gas (which for $\phi=\phi_{\rm FV}$ would happen around $T_{\rm QCD}^{\rm FV}$). However, the dilaton vacuum is non-perturbatively unstable and eventually decays to $\phi_{\rm TV}$ through quantum tunnelling, proceeding through the nucleation of true-vacuum bubbles. The nucleation rate per unit volume is given by $\Gamma/\mathcal{V}=Ae^{-S_{4}}$ [47, 48], with $S_{4}$ the time independent $O(4)$ symmetric bounce action, and $A$ a prefactor. The nucleation temperature, $T_{n}$, is bounded from above by the requirement that the global minimum of $V_{\rm eff}$ is at $\phi_{\rm TV}$. This gives $T_{n}\ll m_{\phi}$ and, more stringently, $T_{n}<T_{c}$, where $T_{c}$ denotes the ‘critical temperature’ at which the two minima have the same energy: $V_{0}(\phi_{\rm TV})+V_{\cal P}(\phi_{\rm TV},T_{c})=V_{0}(\phi_{\rm FV})+V_{\cal P}(\phi_{\rm FV},T_{c})$. Lattice simulations have found that the QCD pressure at strong coupling is lower than the corresponding ideal gas pressure [49, 50] (see [51] for a review). Parametrically, we expect $T_{c}^{3}T_{\rm QCD}\sim\Delta V$ when ${\cal O}(300)\;{\rm MeV}\lesssim T_{c}\lesssim 2\;{\rm GeV}$ and $T_{c}^{4}\sim\Delta V$ for higher critical temperatures, cf. the Supplemental Material [52] for a more detailed discussion. Moreover, the nucleation temperature is bounded from below by the BBN temperature $T_{n}\gtrsim T_{\rm BBN}\approx 4\;{\rm MeV}$ [53, 54, 55, 56]. For $\Delta{\mathfrak{f}}\sim{\cal O}(1)$, $T_{\rm BBN}\gg T_{\rm QCD}^{\rm FV}$, so the false-vacuum QGP does not undergo a QCD cross-over transition prior to the dilaton tunnelling. In the main text of this Letter, we explore $T_{n}$ in the range
$T_{\rm BBN}<T_{n}<T_{\rm QCD}^{\rm TV}\,.$ (8)
The complementary range, $T_{\rm QCD}^{\rm TV}<T_{n}<T_{c}$, naturally predicts ultrarelativistic ‘runaway’ bubbles that, upon collision, generate gravitational wave signals at higher frequencies. Such models can be probed by future gravitational wave detectors, but do not affect the QCD cross-over transition, cf. the Supplemental Material for a more detailed discussion [52]. Fixing the nucleation temperature corresponds to fixing the decay rate through the approximate equation $\Gamma/{\cal V}\simeq H^{4}(T_{n})$, expressing that, at $T_{n}$, one bubble is nucleated per Hubble volume and Hubble time.
@@
Outside the bubbles, the plasma remains in the pre-confined quark-gluon phase (as $T_{n}>T_{\rm QCD}^{\rm FV}$), but inside the bubbles, the temperature is lower than the local QCD confinement scale (i.e. $T_{\rm QCD}^{\rm TV}$), leading to prompt hadronisation and chiral symmetry breaking. In the two quark linear sigma model, the approximate timescale of confinement can be found from $t_{\rm roll}\sim{\cal O}(1/m_{\chi})\ll H^{-1}$, with $m_{\chi}=\sqrt{2\lambda f_{\pi}}$, so chiral-symmetry breaking inside the true-dilaton-vacuum bubbles is approximately instantaneous. Thus, in this scenario, the QCD PT is no longer a smooth cross-over, but characterised by expanding bubbles of the hadronic phase inside the quark-gluon plasma; this scenario effectively turns the QCD phase transition first order. We stress, however, that the mechanism is intrinsically quantum and multiple-field, and lies outside the much-studied framework of thermal QCD PTs.
@@
We require that the transition percolates (at plasma temperature $T_{\star}$) which, for a PT with constant $\Gamma$, requires $\alpha\lesssim 20$ [57], where $\alpha=\Delta V/\rho_{\rm rad}(T_{\rm n})$. For such transitions, the inverse duration of the PT, $\beta$, is set by $\alpha$ and satisfies $8\gtrsim\beta/H_{*}>3$ [57]. As the energy density $\Delta V$ is channelled into the Standard Model, the plasma is reheated to the temperature $T_{\rm rh}=(1+\alpha)^{1/4}T_{n}$, here assuming $T_{\star}\approx T_{n}$. We restrict our discussion to $T_{\rm rh}\lesssim T_{\rm QCD}^{\rm TV}$, so that the hadrons do not deconfine (and then undergo a ‘second’ QCD transition, this time a cross-over).
@@
The subsequent cosmology and observational predictions of this scenario depend on the speed at which the bubbles expand. As we show in the Supplemental Material [52], the bubble walls reach ultrarelativistic terminal velocities before the bubbles collide. This is in notable contrast to BSM or holographic models of thermal first-order QCD PT, in which the latent heat is typically too small to overpower the pressure from the quark-gluon plasma, thereby making ultrarelativistic velocities unattainable.
@@
Gravitational wave signal — As the relativistic dilaton bubbles grow large, collide, and percolate, they produce a stochastic gravitational wave (GW) background, potentially observable today [11, 58]. This is particularly interesting given the recent observation of the Hellings-Down correlation in Pulsar Timing Array (PTA) datasets [4, 5, 6, 7], providing evidence for a stochastic GW background in the nanohertz range. Such a background could potentially be explained astrophysically through supermassive black hole binary mergers [59], but this explanation has difficulties with accommodating the observed amplitude and spectrum [60, 61]. By contrast, several studies have found that GWs generated by a first-order PT provides a better fit to the data [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76] (see [9] for an alternative gravitational wave explanation generated by turbulent sources).
@@
Since the bubbles are expected to reach a terminal velocity before they collide, the main contribution to the GW signal in this scenario is from the dynamics of sound waves generated in the plasma [77, 78, 79, 80, 81] (the contribution to the signal coming from bubble wall collisions, which is relevant for runaway PTs, is discussed in the Supplemental Material [52]).
@@
Thus, the dilaton-induced first-order QCD PT proposed in this work results in peak frequencies in the ${\cal O}(10$$100\,\text{nHz})$ range.
@@
$\times\left(\frac{61.75}{g_{*}(T_{\star})}\right)^{1/3}v_{w}\left(\frac{f}{f_{\rm sw}}\right)^{3}\left(\frac{7}{4+3\left(f/f{\rm sw}\right)^{2}}\right)^{7/2}\;,$ (10)
@@
In Fig. 2 we show the GW energy spectrum generated by the dilaton-induced phase transition for several benchmark values of the parameters. The spectra are overlaid with the posterior distributions from the NANOGrav 15-year free-spectrum analysis [4], shown as violins in two shades of gray (see below). Remarkably, the predictions of our model are broadly compatible with the observed spectrum. In the Supplemental Material [52], we follow [62] and determine the parameter space favoured by the NANOGrav Data by performing a Markov Chain Monte Carlo analysis using the PTArcade code [87, 88]. Only the 14 lowest-frequency bins are used to constrain the spectrum, since the evidence for the GW background is dominated by the low-frequency modes [62]. These bins cover the frequency range $2\;{\rm nHz}\leq f\leq 28\;{\rm nHz}$ and are shaded in darker gray in Fig. 2. For clarity, we also plot the higher-frequency range up to $59\;{\rm nHz}$, shown in lighter gray, that is used for fitting the pulsar-intrinsic red noise. The predicted spectrum provides an excellent fit to the data, with $68\%$ posterior credible intervals given by $11.7\;{\rm MeV}\leq T_{n}\leq 22.7\;{\rm MeV}$ and $0.47\leq\alpha\leq 0.84$. The best-fit spectrum is shown in red in Fig. 2, together with two illustrative examples of spectra for other parameter values. We note that the slight dip in the inferred spectrum above 20 nHz drives the preference for a low nucleation temperature. Higher values of $T_{n}$ push the peak of the spectrum to higher frequencies, while increasing $\alpha$ lowers the peak frequency and increases the overall strength of the signal.
Figure 2: Gravitational wave spectra from a dilaton-induced QCD PT overlaid with NANOGrav 15-year free-spectrum constraints (two shades of gray violins). The red curve corresponds to the best-fit spectrum to the 14 lowest-frequency bins (darker gray). For more information, see the main text.
@@
To realise the hierarchy of Eq. (8) with a nucleation temperature of $T_{n}=100\,{\rm MeV}$ one may, for example, choose $m\sim 10\,{\rm TeV}$, $\Delta\phi\simeq 2\times 10^{-3}~{\rm GeV}$, $\Delta\phi/\Lambda_{1}\sim\mathcal{O}(1)$ and $\Delta V\simeq 0.7\rho_{\rm rad}(T=100\text{MeV})$. In this case, the barrier height in $\phi$ direction is large, and the thin-wall approximation can be used to estimate $S_{4}$ and the decay rate. Thus, our scenario can be realised with a rather minimal set of ingredients.
Discussion — We have shown that quantum tunnelling of the QCD dilaton can turn the QCD phase transition first order and generate a gravitational wave signal intriguingly similar to that observed by Pulsar Timing Arrays. Notably, in our scenario, the PT is induced by multi-field dynamics rather than generated from thermal fluctuations, and its predictions differ quantitatively and qualitatively from the much-studied framework of thermal, first-order QCD transitions.
@@
While the explicit model presented in this work illustrates how it may be realised with rather simple ingredients, it would be interesting to embed this toy model in a more complete theoretical framework.
Finally, the mechanism proposed in this Letter may find applications outside cosmology. Emergent gauge theories with tunable ‘dilaton’ parameters have been discovered in several condensed matter systems [97, 98, 99, 100], and may potentially host non-trivial PTs similar to our scenario. More generally, the essence of our proposal is that the order-parameter space of certain theories can naturally be augmented, leading to rich multiple-field dynamics that can transform the nature of predicted PTs.
Acknowledgments—We thank Mark Hindmarsh, Filippo Sala, and Javier Subils for stimulating discussions. This work was supported by the Swedish Research Council (VR) under grants 2018-03641, 2019-02337 and 2024-04289. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@
In this section, we calculate the pressure exerted on the wall by the quark-gluon plasma and determine the terminal speed of the bubble walls. Recall that we consider percolating phase transitions with time-independent $\Gamma$. The $\beta$-parameter is given by $\beta=\dot{I}(t_{\star})$, with $I(t)$ the average number of bubbles in the past comoving light cone [57], i.e.
$\beta=\frac{4\pi\Gamma}{a(t_{\star})}\int_{0}^{t_{\star}}dt^{\prime}a^{3}(t^{\prime})r^{2}_{\rm com}(t^{\prime},t_{\star})\;,$ (14)
where $a(t)$ denotes the scale factor, $r_{\rm com}(t_{\star},t)$ the comoving bubble radius and $t_{\star}$ the percolation time, defined by $I(t_{\star})=1$. As mentioned in the main text, for transitions that percolate, $0<\alpha\lesssim 20$, which corresponds to $8\gtrsim\beta/H_{\star}>3$
@@
where $g_{q}^{*},\,g^{*}_{g}$ denote, respectively, the effective degrees of freedom for quarks and gluons in the plasma. The inequality assumes $\alpha\gtrsim 3\cdot 10^{-3}\,(\Lambda_{\rm QCD}/T_{n})^{2}$, which holds across the most relevant parameter space. Thus, one-to-one processes alone do not suffice to stop the runaway behaviour [113]  [^4] A more accurate description of friction in a confinement phase transition is given in [114], and results in the scaling $\mathcal{P}_{\rm friction}=\mathcal{P}_{\rm BS}\sim\sum_{i\in BS}\frac{m_{i}^{2}T_{n}^{2}}{24}$, which coincides with that of [113], used here. .
@@
with $\phi_{w}$ a value of the dilaton on the wall and $N_{g}=8$ is the number of gluons. Here, we have assumed that QCD remains perturbative at the wall, i.e. for an intermediate value of $\phi$. This is typically justified since $\Lambda_{\rm QCD}(\phi_{w})\ll\Lambda^{\rm TV}_{\rm QCD}$. The pressure contribution $\mathcal{P}_{1\rightarrow 2}$ grows with $\gamma$ and eventually catches up with $\Delta V$, leading to an ultra-relativistic terminal velocity with gamma factor  [^5] Another potential friction source is bound state formation from string fragmentation [103]. However, due to insignificant supercooling, we expect confinement to happen among quark pairs rather than quarks forming flux-tubes attached to the bubble wall, and string fragmentation to be unimportant. Even in this scenario though, the leading and next-to-leading order friction from gluons has a similar dependence on $\gamma$
@@
Figure S1: Gravitational wave spectra from a dilaton-induced QCD PT overlaid with NANOGrav 15-year free-spectrum constraints (gray violins), including the contribution from bubble wall collisions. Solid lines show the spectra in the regime where the energy is carried by bubble walls, while dashed lines correspond to the regime where the energy is converted into sound waves, for comparison. The solid red curve corresponds to the best fit.
@@
Here $v_{w}=1$ and $\kappa_{\rm coll}$ is the energy fraction that is converted into bubble walls. The corresponding spectra are shown in Fig. S1, where we use several benchmark values of the parameters.